In the following hypothetical reaction A + 3B 
 2C + D, initial moles of A is twice that of B. If
at equilibrium moles of B and C are equal, % of B reacted is
(A) 10% (B) 20% (C) 40% (D) 60%

Grade:11

5 Answers

Havishma Haranath
9 years ago
             A + 3B ----- 2C + D
initially   1     1         0      0
let x moles of B be consumed at equillibrium
equilli    1-x   1-3x    2x     x
given 1-3x=2x
x=1/5
it means 3/5 moles are consumed which means
60 %
Herre B is the limiting reagent
and the intention of taking initial conc. of B as 1 is that initially it is completely unionised.
Havishma Haranath
9 years ago
             A + 3B ----- 2C + D
initially   1     1         0      0
let x moles of B be consumed at equillibrium
equilli    1-x   1-3x    2x     x
given 1-3x=2x
x=1/5
it means 3/5 moles are consumed which means
60 %
Herre B is the limiting reagent
and the intention of taking initial conc. of B as 1 is that initially it is completely unionised.
Havishma Haranath
9 years ago
             A + 3B ----- 2C + D
initially   1     1         0      0
let x moles of B be consumed at equillibrium
equilli    1-x   1-3x    2x     x
given 1-3x=2x
x=1/5
it means 3/5 moles are consumed which means
60 %
Herre B is the limiting reagent
and the intention of taking initial conc. of B as 1 is that initially it is completely unionised.
Havishma Haranath
9 years ago
             A + 3B ----- 2C + D
initially   1     1         0      0
let x moles of B be consumed at equillibrium
equilli    1-x   1-3x    2x     x
given 1-3x=2x
x=1/5
it means 3/5 moles are consumed which means
60 %
Herre B is the limiting reagent
and the intention of taking initial conc. of B as 1 is that initially it is completely unionised.
Havishma Haranath
9 years ago
             A + 3B ----- 2C + D
initially   1     1         0      0
let x moles of B be consumed at equillibrium
equilli    1-x   1-3x    2x     x
given 1-3x=2x
x=1/5
it means 3/5 moles are consumed which means
60 %
Herre B is the limiting reagent
and the intention of taking initial conc. of B as 1 is that initially it is completely unionised.